OpenStudy (anonymous):

help me find the solution set of 3(10/(4-x^2))^2+10/(4-x^2 ) – 10 = 0

OpenStudy (anonymous):

Set $$(4-x^2)=t$$ and solve for t

OpenStudy (anonymous):

It should be a quadratic in t

OpenStudy (anonymous):

assign a certain for 4-x^2 then it will be a quadratic equation

OpenStudy (anonymous):

Yes, then after we solve for t, we know that $$t=4-x^2$$ So we can solve for x

OpenStudy (anonymous):

3t^2+t-10=0

OpenStudy (anonymous):

then factor it?

OpenStudy (anonymous):

number 3

OpenStudy (anonymous):

:) is that the answer?

OpenStudy (anonymous):

the real solution is better :)

OpenStudy (anonymous):

thank you :)